OpenStudy (anonymous):

let f(x)=x2-7x+3 and g(x)=x2-5x+1 find f(x)+g(x) a) x2-8x+7 b) x2-6x+7 c) x3-3x+7 d) x3-7x+7

OpenStudy (anonymous):

Once again, easy....

OpenStudy (anonymous):

All you do is add the two functions together..... f(x) + g(x) = x^2 - 7x +3 + x^2 - 5x + 1 = 2x^2 -12x + 4

OpenStudy (anonymous):

The only thing is, you must have made a mistake because the answers are not matching up

OpenStudy (anonymous):

niti, did you post the question right?